Solve the system using elimination
X+ 3y=5 4x+4y=4

Answer by suruman(21)

x + 3y = 5 - eqn 1
4x + 4y = 4 - eqn 2
Multiplying eqn 1 by 4 on both L.H.S and R.H.S
4x + 3*4y = 5*4
4x + 12y = 20 - eqn 3
(eqn 3) - (eqn 2) =>
8y = 16
y = 16/8 = 2
Substituting value of y in equation 1,
x + 3*2 = 5
x = 5 - (3*2) = 5 - 6 = -1
Thus, x =-1 and y = 2